# On automatic boundedness of operators from ordered to topological vector spaces

### 16 Apr 2026 v3

@@ 1. $i)$
First, we show that if $X_{+}$ is normal then $x_{\alpha}\operatorname{\xrightarrow[]{ru}}0\Longrightarrow x_{\alpha}\operatorname{\xrightarrow[]{o}}0$ in $X$. Suppose $x_{\alpha}\operatorname{\xrightarrow[]{ru}}0$. Then, for some $u\in X_{+}$ there exists an increasing sequence $(\alpha_{n})$ with $\frac{1}{n}u\leq x_{\alpha}\leq\frac{1}{n}u$ for all $\alpha\geq\alpha_{n}$. Let $0\leq y_{0}\leq\frac{1}{n}u$ for all $n\in\mathbb{N}$. Since $\big\|\frac{1}{n}u\big\|\to 0$ then $y_{0}=0$ by [4, Theorem 2.23]. It follows $\frac{1}{n}u\downarrow 0$, and hence $x_{\alpha}\operatorname{\xrightarrow[]{o}}0$.
$i)$ Let $X$ has order continuous norm and $x_{\alpha}\operatorname{\xrightarrow[]{o}}0$ in $X$. Take a net $g_{\beta}\downarrow 0$ in $X$ such that, for each $\beta$ there is $\alpha_{\beta}$ satisfying $\pm x_{\alpha}\leq g_{\beta}$ for $\alpha\geq\alpha_{\beta}$. Find an increasing sequence $(\beta_{k})$ such that $\sup\limits_{z\in[-g_{\beta_{k}},g_{\beta_{k}}]}\|z\|\leq k^{-3}$ for all $k\in\mathbb{N}$, and set $u=\|\cdot\|$-$\sum_{k=1}^{\infty}kg_{\beta_{k}}$. Then $\pm x_{\alpha}\leq g_{\beta_{k}}\leq k^{-1}u$ for $\alpha\geq\alpha_{\beta_{k}}$. Choose any increasing sequence $(\alpha_{k})$ with $\alpha_{k}\geq\alpha_{\beta_{1}},\alpha_{\beta_{2}},\ldots,\alpha_{\beta_{k}}$. Then $\pm x_{\alpha}\leq k^{-1}u$ for $\alpha\geq\alpha_{k}$, and hence $x_{\alpha}\operatorname{\xrightarrow[]{ru}}0$.
ii)ii) The proof is similar. ∎
@@
Let $X$ be an ordered Banach space with a closed generating normal cone and $Y$ a normed space. The following assertions hold.
  1. $i)$
    If $X$ has $\sigma$-order continuous norm then L^σ_Leb(X,Y)=L^σ_onc(X,Y) and L^σ_wLeb(X,Y)=L^σ_owc(X,Y).
  2. ii)ii)
    If $X$ has order continuous norm then L_Leb(X,Y)=L_onc(X,Y) and L_wLeb(X,Y)=L_owc(X,Y).
@@
We prove only $i)$ as the proof of ii)ii) is similar. In view of Lemma 3.9, we need to prove ${\cal L}^{\sigma}_{Leb}(X,Y)\subseteq{\cal L}^{\sigma}_{onc}(X,Y)$ and ${\cal L}^{\sigma}_{wLeb}(X,Y)\subseteq{\cal L}^{\sigma}_{owc}(X,Y)$. Let $x_{n}\operatorname{\xrightarrow[]{o}}0$ in $X$. Then $x_{n}\operatorname{\xrightarrow[]{ru}}0$, by Lemma 3.16, and hence $\|x_{n}\|\to 0$ due to normality of $X_{+}$.
If $T\in{\cal L}^{\sigma}_{Leb}(X,Y)$ then $T\in{\cal L}_{b}(X,Y)$ by Theorem 3.12, and hence $\|Tx_{n}\|\to 0$. If $T\in{\cal L}^{\sigma}_{wLeb}(X,Y)$. Then again $T\in{\cal L}_{b}(X,Y)$ by Theorem 3.12. So $\|Tx_{n}\|\to 0$ and $Tx_{n}\operatorname{\xrightarrow[]{w}}0$. ∎
@@
Let $E$ be a Banach lattice with order continuous norm and $Y$ a normed space. Then ${\cal L}^{\sigma}_{Leb}(E,Y)={\cal L}^{\sigma}_{onc}(E,Y)$, ${\cal L}_{Leb}(E,Y)={\cal L}_{onc}(E,Y)$, ${\cal L}^{\sigma}_{wLeb}(E,Y)={\cal L}^{\sigma}_{owc}(E,Y)$, and ${\cal L}_{wLeb}(E,Y)={\cal L}_{owc}(E,Y)$.
Finally, we give some conditions on order bounded operators to be order-to-norm continuous.
@@
Let $X$ be an ordered Banach space with a closed normal cone and $Y$ an ordered normed space with a generating normal cone. The following assertions hold.
  1. $i)$
    If $X$ has $\sigma$-order continuous norm then ${\cal L}_{ob}(X,Y)\subseteq{\cal L}^{\sigma}_{onc}(X,Y)$.
  2. ii)ii)
    If $X$ has order continuous norm then ${\cal L}_{ob}(X,Y)\subseteq{\cal L}_{onc}(X,Y)$.
@@
We prove only $i)$ as the proof of ii)ii) is similar. Let $T\in{\cal L}_{ob}(X,Y)$ and $x_{n}\operatorname{\xrightarrow[]{o}}0$ in $X$. By Lemma 3.16, $x_{n}\operatorname{\xrightarrow[]{ru}}0$. It follows from [15, Corollary 2.5] that $Tx_{n}\operatorname{\xrightarrow[]{ru}}0$. The normality of the cone $Y_{+}$ ensures $\|Tx_{n}\|\to 0$. ∎

### 20 Apr 2026 v4

@@ 3. $d)$
Since $X_{+}$ is normal, $X$ is Archimedean, and hence $x_{\alpha}{\xrightarrow[]{ru}}0\Longrightarrow x_{\alpha}{\xrightarrow[]{o}}0$ in $X$.
Let $X$ has order continuous norm and $x_{\alpha}{\xrightarrow[]{o}}0$ in $X$. Take a net $g_{\beta}\downarrow 0$ in $X$ such that, for each $\beta$ there is $\alpha_{\beta}$ satisfying $\pm x_{\alpha}\leq g_{\beta}$ for $\alpha\geq\alpha_{\beta}$. Since the norm in $X$ is order continuous and $X_{+}$ is normal, there exists an increasing sequence $(\beta_{k})$ such that $\sup\limits_{z\in[-g_{\beta_{k}},g_{\beta_{k}}]}\|z\|\leq k^{-3}$ for all $k\in\mathbb{N}$. Set $u=\|\cdot\|$-$\sum_{k=1}^{\infty}kg_{\beta_{k}}$. Then $\pm x_{\alpha}\leq g_{\beta_{k}}\leq k^{-1}u$ for $\alpha\geq\alpha_{\beta_{k}}$. Choose any increasing sequence $(\alpha_{k})$ with $\alpha_{k}\geq\alpha_{\beta_{1}},\alpha_{\beta_{2}},\ldots,\alpha_{\beta_{k}}$. Then $\pm x_{\alpha}\leq k^{-1}u$ for $\alpha\geq\alpha_{k}$, and hence $x_{\alpha}{\xrightarrow[]{ru}}0$. ∎
@@
Let $X$ be an ordered Banach space with a closed generating normal cone and $Y$ a normed space. If $X$ has order continuous norm then
${\cal L}_{Leb}(X,Y)={\cal L}_{otc}(X,Y)\ \ and\ \ {\cal L}_{wLeb}(X,Y)={\cal L}_{owc}(X,Y).$
@@
In view of Lemma 3.8, we need to prove ${\cal L}_{Leb}(X,Y)\subseteq{\cal L}_{otc}(X,Y)$ and ${\cal L}_{wLeb}(X,Y)\subseteq{\cal L}_{owc}(X,Y)$. Let $x_{\alpha}{\xrightarrow[]{o}}0$ in $X$. Then $x_{\alpha}{\xrightarrow[]{ru}}0$ by Lemma 3.15, and hence $\|x_{\alpha}\|\to 0$ due to normality of $X_{+}$.
If $T\in{\cal L}_{Leb}(X,Y)$ then $T\in{\cal L}^{\sigma}_{Leb}(X,Y)$. By Theorem 3.12, $T\in{\cal L}_{b}(X,Y)$, and hence $\|Tx_{\alpha}\|\to 0$. Thus, $T\in{\cal L}_{otc}(X,Y)$. If $T\in{\cal L}_{wLeb}(X,Y)$ then again $T\in{\cal L}_{b}(X,Y)$ by Theorem 3.12. Thus, $\|Tx_{\alpha}\|\to 0$, and hence $Tx_{\alpha}{\xrightarrow[]{w}}0$. So, $T\in{\cal L}_{owc}(X,Y)$. ∎
@@
Let $E$ be a Banach lattice with order continuous norm and $Y$ a normed space. Then, ${\cal L}_{Leb}(E,Y)={\cal L}_{otc}(E,Y)$ and ${\cal L}_{wLeb}(E,Y)={\cal L}_{owc}(E,Y)$.
Finally, we give the conditions for order-to-topology continuity of order bounded operators.
@@
Let $X$ be an ordered Banach space with a closed normal cone and order continuous norm, $Y$ be an ordered normed space with a generating normal cone, and $T:X\to Y$ and order bounded operator. Then $T$ is order-to-topology continuous.
@@
Let $x_{\alpha}{\xrightarrow[]{o}}0$ in $X$. By Lemma 3.15, $x_{\alpha}{\xrightarrow[]{ru}}0$. It follows from [15, Corollary 2.5] that $Tx_{\alpha}{\xrightarrow[]{ru}}0$. The normality of the cone $Y_{+}$ ensures $\|Tx_{\alpha}\|\to 0$. ∎